OpenStudy (anonymous):

Write the equation for a circle with its center at (2,-5) and a radius of 6.

OpenStudy (anonymous):

write \[(x-h)^2+(y-k)^2=r^2\] with \[h=2, k=-5, r =6\]

OpenStudy (anonymous):

The Answer Choices Are: A. (x + 2)2 + (y - 5)2 = 6 B. (x + 2)2 + (y - 5)2 = 36 C. (x - 2)2 + (y + 5)2 = 6 D. (x - 2)2 + (y + 5)2 = 36

OpenStudy (anonymous):

lhm